# Are Statistical Methods Obsolete in the Era of Deep Learning?

### v1

@@ 4.1 Experimental setup
We use the same setup of the PINN and MAGI as in Section 3.1 for consistent comparison. The PINN hyperparameter $\lambda$ that governs the balance between the $L_{2}$ and the physics loss (see Equation (3)) will be enumerated at $\lambda=0.1,1,10,100,$ and $1000$ for cross-board comparison. For all experiments in this section, the hyperparameters in MAGI are left to be automatically fitted. Other configuration details regarding PINN and MAGI are analogous to those of the fully-observed case in Section 3.1 and are included in the SI Section S1 and Section S2 respectively.
@@ 4.2 Results from chaotic (butterfly) regime
Figure 8: MAGI and PINN reconstructed trajectories on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. Panel layouts are the same as Figure 1.
Visually from Figure 8, PINN struggles to fit the data, particularly at $\lambda=10$, where PINN tends to produce a flat-line prediction for the estimated trajectory, attributing all the deviations from the flat line to observation noise. This pattern persists over a broader range of $\lambda$ values, as demonstrated in SI Figure S4. Even when the observation error penalty $\lambda$ is increased to 1000, this behavior remains noticeable. At $\lambda=1000$, while some curves still underfit, other curves start to overfit the noise, suggesting an inability to balance the trade-off between flat-line fitting and overfitting. To rule out the possibility of optimization issues, we attempted a warm-start for PINN by starting without the physics loss, allowing the neural network to solely interpolate the data. Yet, even after this warm start, the flat prediction re-emerged after optimization with physics loss, indicating that this behavior is inherent to the NN model design rather than merely an optimization limitation. Mathematically, this flat-line output behavior can be explained by the fact that a constant trajectory solution trivially satisfies the governing ODEs of the Lorenz system. Moreover, an over-parameterized neural network with random initialization is prone to producing a flat line output at initialization. Gradient-based optimization further exacerbates this issue by naturally inducing shrinkage toward the flat-line initialization, compounding the problem (Glorot and Bengio, 2010).
On the other hand, MAGI’s trajectory estimates, as shown in Figure 8, exhibit both low bias and low variance. Additionally, the variance across trials remains relatively uniform over time, regardless of the presence of peaks or troughs in the data. Quantifying these visual observations with the RMSE of the in-sample trajectory reconstruction, as displayed in the boxplots in Figure 9, reveals that MAGI consistently outperforms all tested PINN variants. MAGI’s superior performance may be attributable to the fact that GP priors favor oscillatory curves over flat-line solutions.
@@
Figure 9: Boxplots showing trajectory reconstruction RMSEs across $100$ datasets on the Lorenz Chaotic (Butterfly) case. Lower values indicate better performance. The three panels correspond to the three system components $X,Y$, and $Z$. Panel layouts are the same as Figure 2.
@@
For parameter inference, Figure 10 presents the boxplot of absolute estimation errors. While MAGI maintains a clear advantage over PINN, certain PINN configurations can still provide reasonable parameter estimates, as observed in the previous SEIR example in Section 3. Notably, when $\lambda=100$, the PINN parameter estimates are relatively accurate on $\sigma$. This can be explained by the nature of PINN’s fitted trajectory: although flat-line segments provide no information about the ODE parameters, the periods where PINN successfully captures the ups and downs of the data reveal meaningful parameter information. As a result, the PINN estimates of parameters are not as poor as the trajectory RMSE might suggest, although still not as efficient as MAGI. This observation aligns with the SEIR case, demonstrating that PINN, much like a two-stage method, does not require a perfect trajectory estimate to produce reasonable parameter estimates. However, hyperparameter tuning in PINN is still necessary: only $\lambda=100$ achieves the best trade-off between flat-line output and overfitting the observation noise. In contrast, MAGI requires no hyperparameter tuning at all in any of these Lorenz experiments.
@@
Figure 10: Boxplots showing parameter inference absolute errors across $100$ datasets on the Lorenz Chaotic (Butterfly) testbed. Lower values indicate better performance. Intra-panel layouts are identical to Figure 9.
Finally, Figure 11 illustrates the forecasting performances of MAGI and PINN on the Chaotic (Butterfly) case for $\lambda=10$ and $\lambda=1000$ where PINN’s tendency of flat-line prediction is observed again during the forecasting period. In this low noise case, PINN’s in-sample behavior is quite strong, yet it still struggles with the forward prediction across $\lambda$ values (SI Figure S5), which is consistent with our observation from the SEIR example. MAGI can consistently and accurately forecast the future prediction interval up to time $t=5$, where it starts to reach the limit on this Chaotic (Butterfly) regime.
@@
Figure 11: MAGI and PINN forecasted trajectories on the Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The panel layouts are the same as Figure 1.
@@
It could be argued that the well-known Chaotic (Butterfly) regime of the Lorenz system is notoriously sensitive to parameter variations, making its inverse problem potentially easier. A small change in parameters can lead to significant changes in the system trajectories, implying that observing the trajectories provides strong information about the parameters. To further investigate PINN and MAGI, we thus conducted a second set of experiments under a stable Lorenz regime, where we observed similar behavior in PINN and MAGI as in the chaotic regime. The flat-line output of PINN remains evident across a wide range of $\lambda$ values. MAGI continues to demonstrate consistently superior performance across datasets and time periods. Detailed analysis of this stable regime can be found in SI Section S5.1.
5 Discussion
@@
This study highlights the relevance and potential of statistically principled methods like MAGI in addressing complex systems, even in the era of deep learning. From our analyses on the SEIR and Lorenz systems, it is evident that MAGI outperforms PINN, particularly in more complex scenarios such as those involving missing components and/or chaotic behavior. Furthermore, as a Bayesian method, MAGI offers uncertainty quantification through credible intervals, a capability that PINN inherently lacks. For instance, on the SEIR model, SI Table S1 shows the frequentist coverage of the 95% credible intervals by MAGI, where the actual coverage is reasonably close to the nominal 95% level for all system parameters, both in the fully observed case and the missing $E$ component case.
@@
It is also of interest to investigate how close the reconstructed solutions of the ODEs follow the governing mechanistic equations. To quantify this, we inspect the values of the mechanistic fidelity loss component of PINN in Equation (3) and MAGI’s counterpart in Equation (7), using the standard Euclidean norm, instead of over $C^{-1}$ for a meaningful comparison. The results for the Lorenz Chaotic (Butterfly) case are shown in Figure S1, where we observe that MAGI shows considerably better mechanistic fidelity with respect to the original ODE governing equations than PINN variants with $\lambda\geq 10$, which are the better-performing PINN configurations for inference. PINNs with $\lambda=0.1$ or $1.0$ have similar level of mechanistic fidelity to MAGI but they sacrifice significant inference performance. Since a flatlined solution trivially satisfies the Lorenz ODEs, mechanistic fidelity is only relevant and useful in combination with strong inference performance. MAGI’s higher mechanistic fidelity may also contribute to its ability to generalize with future predictions.
Statistical methods like MAGI, derived from a full ground-up probabilistic modeling of the data generating mechanism, can often avoid the need for extensive regularization and ad hoc modifications, as typically required by the training of deep learning methods. For example, MAGI is capable of automatic hyperparameter setting in most cases. We emphasize that, despite being a Bayesian method, MAGI’s prior distributions on its parameters do not require manual domain-specific specification: almost everything is fully-automated.
@@
Now, returning to the question posed in the article title, we argue that statistical methods are far from obsolete. Instead, they provide complementary strengths to neural network approaches, particularly in scenarios with limited amount of data and in scenarios that demand interpretability, robustness, and reliable uncertainty quantification. We believe future research that integrates the complementary strengths of statistical methods like MAGI with neural network innovations holds the promise of significantly advancing the field of computational modeling in the era of AI.
@@
Detailed implementation code at https://anonymous.4open.science/r/stat-vs-dl-71B2/. (anonymous GitHub code repository)
@@
The PINN is implemented as a fully-connected neural network, also referred to as a multi-layer perceptron (MLP), using TensorFlow 2.17.0 (Abadi et al., 2015). The architecture consists of 3-4 hidden layers, each with a width of 20 neurons. The activation function for all layers is the hyperbolic tangent ($\tanh$), chosen for its smoothness and suitability for modeling continuous physical processes. Our implementation is gratefully borrowed from van Herten et al. (2022)’s formulation of PINN for the Lorenz system, replicating Lu et al. (2021)’s DeepXDE library.
The network is trained using the Adam optimizer, with a fixed learning rate set to $0.01$. The training process spans $60000$ epochs to ensure convergence and adequate fitting of the data to the governing physical laws, as demonstrated in van Herten et al. (2022). The physics loss is evaluated at the same discretized time points used by the MAGI method, as described in the main text. This ensures consistency in the temporal resolution between the two comparison methods.
The full implementation details and codebase for the PINN are available at the following anonymized GitHub repository: https://anonymous.4open.science/r/stat-vs-dl-71B2.
@@
For the implementation of MAGI, we utilize a Gaussian process (GP) kernel based on the Matérn kernel with 2.01 degrees of freedom. This kernel choice balances flexibility and smoothness, making it well-suited for modeling the dynamics of the system. The implementation is built using TensorFlow Probability 0.24.0 (Dillon et al. (2017)), leveraging its built-in advanced Markov Chain Monte Carlo capabilities.
The hyperparameters of the Gaussian process are estimated through GP smoothing of the observed data, as previously documented in Yang et al. (2021). For missing components of the system, the hyperparameters are determined through one of two strategies: (1) setting them to fixed values; or (2) employing a novel interpolation and optimization routine discussed in the following subsection. In general, we perform $3000$ burn-in steps followed by $3000$ sampling steps, with full code in our anonymized GitHub repository. Please see Section S2.2 for forecasting details.
@@
    First, we upgraded the base Hamiltonian Monte Carlo (Hamiltonian Monte Carlo Neal (2011)) (HMC) sampler from Yang et al. (2021) to the No U-Turn Sampler (Hoffman and Gelman (2014)), itself a variant of HMC, with Dual-Averaging Stepsize Adaptation (Nesterov (2009)) to enable a completely tuning-free setup, leveraging TensorFlow Probability’s collection of advanced Monte Carlo samplers. Our implementation is fully-compatible with GPUs and XLA integration for maximum computational power. For future-proofing, we include a toggleable option for logarithmic annealing of the log-posterior function to encourage additional exploration of the state space, but this is not enabled in any of our experiments.
@@
    Fifth, we introduce a novel procedure for interpolating missing components’ observations and fitting corresponding kernel hyperparameters. First, we use a second-order gradient-matching procedure optimized via Adam to jointly estimate the missing components’ trajectory values and the initial values for $\bm{\Theta}$. Second, we apply GP smoothing on the interpolated missing components’ trajectory values to estimate the kernel hyperparameters. This is a departure from the less-stable method used in Yang et al. (2021) that directly tries to maximize the MAGI log-posterior function, and in the spirit of modern “differentiable almost everything” paradigms in machine learning (see ICML 2024 Workshop on Differentiable Almost Everything).
@@
    Finally, to provide smoother initializations of $\mathbf{\hat{X}}(t)$ for the NUTS sampler and reduce the risk of the sampler becoming stuck in local modes, we introduce a cross-validated cubic-spline-based smoothing mechanism to reduce jaggedness in the original linearly-interpolated initializations of $\mathbf{\hat{X}}(t)$
The complete codebase for the improved MAGI algorithm is available at the following anonymized GitHub repository: https://anonymous.4open.science/r/stat-vs-dl-71B2.
@@
Figure S2: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the fully observed case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Each solid blue curve is the estimate from one dataset. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
@@
Figure S3: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the partially-observed case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Each solid blue curve is the estimate from one dataset. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
S5 Full Lorenz chaotic (butterfly) trajectory figures

### v2

@@ 4.1 Experimental setup
We follow very similar PINN and MAGI implementation protocols as in the SEIR experiments (Section 3.1) to ensure a consistent comparison across systems. For PINN, we enumerate the balancing hyperparameter $\lambda$ in $\lambda=0.1,1,10,100,$ and $1000$ (Equation (3)). We evaluate a broad set of PINN configurations by varying the implementation: a practitioner-style codebase used in the PINN literature (van Herten et al. 2022) versus the DeepXDE package (Lu et al. 2021, Wang et al. 2022), network size, learning-rate scheduling, and optimizer routine. As in Section 3.1, we select the best-performing configuration per $\lambda$ value based on inverse-problem accuracy for in-sample experiments, or forecast accuracy for forecasting problems, and only report these best-case PINN results in the main text; comprehensive comparisons across the remaining PINN configurations are deferred to the SI. In the remainder of this Lorenz section, we refer to the best-case PINN simply as PINN whenever it is clear from context.
For MAGI, for all Lorenz experiments, hyperparameters are automatically estimated by marginal likelihood maximization (Section 3.1.2) and are not manually tuned. For in-sample experiments, we perform $3,000$ burn-in and $3,000$ sampling steps. For forecasting, because of the chaotic nature of the Lorenz system, we use the sequential forecasting setup described in SI Section S2 with $3,000$ burn-in and $3,000$ sampling iterations in the initial fit and per step. Remaining configuration details are summarized in SI Sections S1 and S2.
@@ 4.2 Results from chaotic (butterfly) regime
Figure 8: MAGI and best-case PINN reconstructed trajectories on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. Panel layouts are the same as Figure 1.
Visually Figure 8 shows that MAGI reconstructs the Lorenz trajectories with both low bias and low trial-to-trial variability: across the $100$ replicated datasets, the reconstructed curves (blue) closely hug the ground-truth trajectory (black dashed) for all three coordinates $(X,Y,Z)$. Moreover, MAGI’s uncertainty appears fairly uniform over time: the spread of reconstructions does not noticeably inflate around peaks or troughs.
In contrast, even the best-performing PINN settings shown above (selected post-hoc: $\lambda=10,100,1000$) exhibit substantially larger variability across trials: reconstructed trajectories look visibly “fuzzier,” especially during the initial transient and oscillatory segments. This is likely because these PINNs are overfitting to the noisy data points. Looking across the broader $\lambda$-sweep in Figure S8, smaller $\lambda$ values perform markedly worse: the learned trajectories tend to collapse towards near constant (flat) solutions, effectively attributing most deviations from the flat line to observation noise.
Mathematically, this flat-line output behavior can be explained by the fact that a constant trajectory solution trivially satisfies the governing ODEs of the Lorenz system. Moreover, an over-parameterized neural network with random initialization is prone to producing a flat line output at initialization (especially for small $\lambda$, i.e. underweighting the data-fit term). Gradient-based optimization further exacerbates this issue by naturally inducing shrinkage toward the flat-line initialization, compounding the problem (Glorot and Bengio 2010).
Quantitatively, the RMSE boxplots in Figure 9 corroborate these visual patterns: MAGI achieves lower in-sample reconstruction error than every PINN variant we tested. One plausible explanation is that MAGI’s GP priors and principled uncertainty modeling discourage degenerate flat-line solutions and thus favor oscillatory curves that better match the underlying dynamics.
@@
Figure 9: Boxplots showing trajectory reconstruction RMSEs across $100$ datasets on the Lorenz Chaotic (Butterfly) case. Lower values indicate better performance. The three panels correspond to the three system components $X,Y$, and $Z$. Panel layouts are the same as Figure 2. Only best-case PINN configurations are reported here. Results for other inferior PINN configurations are in SI Figure S10.
For parameter inference, Figure 10 presents boxplots of absolute estimation errors for $(\beta,\rho,\sigma)$. For $\beta$, MAGI tends to have the smallest errors overall compared to all tested PINN settings. For $\rho$, MAGI is broadly comparable to the best post-hoc selected PINN variants ($\lambda\in\{10,100,1000\}$), while small-$\lambda$ PINNs perform dramatically worse. Finally, for $\sigma$, certain intermediate-$\lambda$ PINNs (notably, $\lambda=10$ and $\lambda=100$) can attain smaller errors than MAGI in this experiment, while overly small or overly large $\lambda$ values degrade performance considerably. Overall, the key takeaway is that PINN performance remains highly $\lambda$-sensitive: underweighting the data-fit term yields large errors consistent with trajectory-collapse behavior, while excessively aggressive data-fitting can impede parameter recovery.
@@
One plausible explanation for PINN’s relatively strong performance on $\sigma$ is that while flat-line segments provide little information about the ODE parameters, the periods where PINN successfully captures the oscillatory structure can still reveal meaningful parameter information. As a result, PINN’s parameter estimates are not as poor as its trajectory RMSE might suggest. This observation aligns with the SEIR case, demonstrating that PINN, much like a two-stage method, does not require a perfect trajectory estimate to produce reasonable parameter estimates.
Nonetheless, we emphasize that MAGI delivers strong performance without any manual hyperparameter tuning across these Lorenz settings, while PINN requires extensive hyperparameter selection to obtain reasonably comparable results.
@@
Figure 10: Boxplots showing parameter inference absolute errors across $100$ datasets on the Lorenz Chaotic (Butterfly) testbed. Lower values indicate better performance. Intra-panel layouts are identical to Figure 9. Only best-case PINN configurations are reported here. Results for other inferior PINN configurations are in SI Figure S11.
Finally, Figure 11 compares the forecasting behavior of MAGI and PINN on the Chaotic (Butterfly) regime. MAGI continues to tightly track the ground-truth trajectory closely through the forecasting window, with relatively little inter-trial dispersion in all three coordinates. In contrast, PINN forecasts are much less stable across datasets and highly sensitive to $\lambda$. For $\lambda=10$, many PINN forecasts rapidly diverge after the end of the in-sample region and often produce qualitative incorrect forecast trajectories (including collapse towards near-constant behavior in some trials). For $\lambda=100$, all PINN forecasts appears to collapse towards near-constant behavior. For $\lambda=1000$, PINN does not collapse as severely, but its forecast spread grows substantially over time, indicative of increased sensitivity/instability in out-of-sample dynamics.
Overall, even in this low-noise setting where PINN can fit the in-sample segment quite well, it struggles to produce reliable forward predictions across $\lambda$ values (see also SI Figure S9), consistent with our observation from the SEIR example. In contrast, MAGI can consistently and accurately forecast the future prediction interval up to time $t=5$: beyond this range, the intrinsic sensitivity of chaotic dynamics is expected to impede long-horizon predictability for any numerical method.
@@
Figure 11: MAGI and best-case PINN forecasted trajectories on the Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The panel layouts are the same as Figure 1.
@@
It could be argued that the well-known Chaotic (Butterfly) regime of the Lorenz system is notoriously sensitive to parameter variations, making its inverse problem potentially easier. A small change in parameters can lead to significant changes in the system trajectories, implying that observing the trajectories provides strong information about the parameters. To further investigate PINN and MAGI, we thus conducted a second set of experiments under a stable Lorenz regime, where MAGI and PINN performance dynamics are similar to those in the chaotic regime. Detailed analysis of this stable regime can be found in SI Section S6.
5 Discussion and Conclusion
@@
This study highlights the relevance and potential of statistically principled methods like MAGI in addressing complex systems, even in the era of deep learning. From our analyses on the SEIR and Lorenz systems, it is evident that MAGI routinely outperforms PINN, particularly in more complex scenarios such as those involving missing components and/or chaotic behavior, despite having orders-of-magnitude fewer parameters. Furthermore, as a Bayesian method, MAGI offers uncertainty quantification through credible intervals, a capability that PINN inherently lacks. For instance, on the SEIR model, SI Table S1 shows the frequentist coverage of the 95% credible intervals by MAGI, where the actual coverage is reasonably close to the nominal 95% level for all system parameters, both in the fully observed case and the missing $E$ component case.
@@
It is also of interest to investigate how close the reconstructed solutions of the ODEs follow the governing mechanistic equations. To quantify this, we inspect the values of the mechanistic fidelity loss component of PINN in Equation (3) and MAGI’s counterpart in Equation (7), using the standard Euclidean norm (instead of over $C^{-1}$) for a meaningful comparison. The results for the Lorenz Chaotic (Butterfly) case are shown in Figure S1, where we observe that MAGI shows considerably better mechanistic fidelity with respect to the original ODE governing equations than PINN variants with $\lambda>10$, which are the better-performing PINN configurations for trajectory inference. PINNs with $\lambda=0.1$ or $1.0$ have slightly better or similar levels of mechanistic fidelity to MAGI, but they sacrifice significant inference performance. Since a flatlined solution trivially satisfies the Lorenz ODEs, mechanistic fidelity is only relevant and useful in combination with strong trajectory inference performance. Nonetheless, MAGI’s high mechanistic fidelity—and its apparent robustness to flat-line trajectory collapse—may also contribute to its ability to generalize with future predictions.
Statistical methods like MAGI, derived from a full ground-up probabilistic modeling of the data generating mechanism, can often avoid the need for extensive regularization and ad hoc modifications, as typically required by the training of deep learning methods. For example, MAGI is capable of automatic hyperparameter setting in most cases. On the other hand, PINN can be sensitive to the choice of loss function weighting (Dai et al. 2024). Therefore, we intentionally tested a wide range of configurations for PINN by investigating different choices of (a) the weighting parameter $\lambda$; (b) neural network architecture; (c) learning rate scheduler; and (d) optimizer combinations, and reported the best performing variants per $\lambda$. Despite this effort, PINN oftentimes failed to capture the physics in the Lorenz system. While the sensitivity of PINN to the depth and width of the network typically stabilizes once the network becomes sufficiently large (Krishnapriyan et al. 2021), there are also reported sensitivities to other aspects of the DNN structure, such as the activation function (Wang et al. 2023a), collocation point sampling (Wu et al. 2023), and other training techniques (Cho et al. 2024, Wang et al. 2024). Nevertheless, it is evident that PINN requires significantly more careful tuning than MAGI.
The contrast between MAGI and PINN can also be viewed through the lens of what an analyst must specify and what they obtain in return. Bayesian GP-based formulations like MAGI require the analyst to make modeling assumptions explicit (e.g., observation noise, smoothness, and how the differential equation constrains the latent trajectory), but these assumptions are rewarded with coherent uncertainty quantification for parameters, trajectories, and predictions. In contrast, standard PINNs encode the governing equation and data primarily through a penalized optimization objective and are often used as general-purpose tools, but their performance can depend strongly on optimization and tuning choices mentioned above, and they do not natively deliver calibrated posterior uncertainty. More broadly, performance gains arise from exploiting problem structure rather than from universally applicable optimization alone, consistent with the “No Free Lunch” perspective that no method can dominate across all problems without leveraging problem-specific assumptions (Wolpert and Macready 1997).
@@
Now, returning to the question posed in the article title, we argue that statistical methods are still very relevant for inference problems, even in the age of deep-learning and AI. Instead, they provide complementary strengths to neural network approaches, particularly in scenarios with limited amount of data and in scenarios that demand interpretability, robustness, and reliable uncertainty quantification. Our findings in this study are consistent with the long-standing belief that classical statistical or mechanistic models typically outperform neural networks when the model structure is well-specified and accurately reflects the underlying data-generating process. Nevertheless, despite these advantages, there has been a noticeable surge in the popularity of PINN in recent literature, often at the expense of more classical statistical or Bayesian methods such as MAGI, even in cases where the limitations of PINNs are evident. We believe future research that integrates the complementary strengths of statistical methods like MAGI with neural network innovations holds the promise of significantly advancing the field of computational modeling in the era of AI.
Disclosure statement
The authors declare no conflicts of interest.
Data Availability Statement
Simulated data and detailed implementation code have been made available at https://anonymous.4open.science/r/tas-revisions-7650/. (anonymous GitHub code repository)
@@
Detailed implementation code at https://anonymous.4open.science/r/tas-revisions-7650/. (anonymous GitHub code repository)
@@
We explored two versions of PINN, both implemented as fully-connected neural networks: (a) a practitioner-style codebase based on van Herten et al. (2022) implemented with TensorFlow 2.17.0 (Abadi et al. 2015); and (b) a framework using the widely-adopted DeepXDE framework Lu et al. (2021) with a PyTorch 2.10 backend (Paszke et al. 2019).
S1.1 Practitioner-style codebase based on van Herten et al. (2022)
The PINN is implemented as a fully-connected neural network, also referred to as a multi-layer perceptron (MLP). For Lorenz experiments, we use 3 hidden layers with 32 neurons per layer, with batch normalization. For SEIR experiments, we use 5 hidden layers with 32 neurons per layer, with batch normalization. In both cases, the networks are trained using the Adam optimizer with a fixed learning rate of $0.01$ for $60,000$ epochs. The activation function for all layers is the hyperbolic tangent ($\tanh$), chosen for its smoothness and suitability for modeling continuous physical processes. We sweep over 5 values of the $L_{2}$-vs.-physics-loss governor hyperparameter $\lambda\in\{0.1,1.0,10.0,100.0,1000.0\}$. ODE parameters for both Lorenz and SEIR are modeled on the log-scale. ODE components for Lorenz are modeled on the original scale, while ODE components for SEIR are modeled on the log-scale for numerical stability.
Our implementation is gratefully borrowed from van Herten et al. (2022)’s formulation of PINN for the Lorenz system.
S1.2 DeepXDE-based framework based on Lu et al. (2021)
The PINN is also implemented as a fully-connected neural network. We sweep over 5 values of the $L_{2}$-vs.-physics-loss governor hyperparameter $\lambda\in\{0.1,1.0,10.0,100.0,1000.0\}$. We also sweep over (i) 2 network layer-size configurations (3 hidden layers of 40 or 512 units); (ii) 2 learning rate scheduler choices for initial first-order Adam optimization (constant at $0.001$ or exponentially decaying by $0.9$ every $5,000$ epochs with initial learning rate $0.001$); (iii) whether to fine-tune with second-order L-BFGS after first-order Adam (Rathore et al. 2024). If a fixed learning rate is used, we run Adam for $60,000$ epochs; if the exponentially-decaying learning rate is used, we run Adam for $300,000$ epochs, following the best practices demonstrated in Lu et al. (2021) and Wang et al. (2024). ODE parameters for both Lorenz and SEIR are modeled on the log-scale. ODE components for Lorenz are modeled on the original scale, while ODE components for SEIR are modeled on the log-scale for numerical stability. We do note that the fully-connected neural network for SEIR outputs with a log-softmax transformation to ensure that $S+E+I+R=1$ with $S,E,I,R\geq 0$: this is necessary to ensure numerical stability of DeepXDE.
To emphasize, we select the best (i.e., lowest parameter error for in-sample problems or forecasting error for forecasting problems) PINN hyperparameter combination per $\lambda$ over 5 seeds to make the comparison conservative in PINN’s favor and then re-run the selected configurations for a total of 100 seeds.
The full implementation details and codebases for the PINNs are available at the following anonymized GitHub repository: https://anonymous.4open.science/r/tas-revisions-7650.
@@
For the implementation of MAGI, we utilize a Gaussian process (GP) kernel based on the Matérn kernel with 2.01 degrees of freedom. This kernel choice balances flexibility and smoothness, making it well-suited for modeling the dynamics of the system. The implementation is built using TensorFlow Probability 0.25.0 (Dillon et al. (2017)), leveraging its built-in advanced Markov Chain Monte Carlo capabilities.
The hyperparameters of the Gaussian process are estimated through GP smoothing of the observed data, as previously documented in Yang et al. (2021). For missing components of the system, the hyperparameters are determined through one of two strategies: (1) setting them to fixed values; or (2) employing a novel interpolation and optimization routine discussed in the following subsection. Please see Section S2.2 for forecasting details.
For Lorenz, ODE components are sampled on the original scale, while for SEIR, they are sampled on the log-scale. ODE parameters for both the Lorenz and SEIR testbeds are sampled on the original scale. Numbers of MCMC burn-in and sampling steps are described in Sections 3.1.2 and 4.1. For Lorenz (Chaotic Butterfly), Lorenz (Transient Chaos), and SEIR (fully observed components), we use MAGI’s automatic hyperparameter tuning routines (no user-side intervention needed). We only manually specify hyperparametres for the SEIR (missing component) setting.
@@
The full implementation details and codebase for MAGI are available at the following anonymized GitHub repository: https://anonymous.4open.science/r/tas-revisions-7650
@@
    First, we upgraded the base Hamiltonian Monte Carlo (Hamiltonian Monte Carlo (Neal 2011)) (HMC) sampler from Yang et al. (2021) to the No U-Turn Sampler (Hoffman and Gelman (2014)), itself a variant of HMC, with Dual-Averaging Stepsize Adaptation (Nesterov (2009)) to enable a completely tuning-free setup, leveraging TensorFlow Probability’s collection of advanced Monte Carlo samplers. Our implementation is fully-compatible with GPUs and XLA integration for maximum computational power. For future-proofing, we include a toggleable option for logarithmic annealing of the log-posterior function to encourage additional exploration of the state space, but this is not enabled in any of our experiments.
@@
    Fifth, we introduce a novel procedure for interpolating missing components’ observations and fitting corresponding kernel hyperparameters, all without the need for manual user intervention. First, we use a second-order gradient-matching procedure optimized via Adam to jointly estimate the missing components’ trajectory values and the initial values for $\bm{\Theta}$. Second, we apply GP smoothing on the interpolated missing components’ trajectory values to estimate the kernel hyperparameters. This is a departure from the less-stable method used in Yang et al. (2021) that directly tries to maximize the MAGI log-posterior function, and in the spirit of modern “differentiable almost everything” paradigms in machine learning (see ICML 2024 Workshop on Differentiable Almost Everything).
@@
    Finally, to provide smoother initializations of $\mathbf{\hat{X}}(t)$ for the NUTS sampler and reduce the risk of the sampler becoming stuck in local modes, we introduce a cross-validated cubic-spline-based smoothing mechanism to reduce jaggedness in the original linearly-interpolated initializations of $\mathbf{\hat{X}}(t)$, which also does not require manual user intervention.
The complete codebase for the improved MAGI algorithm is available at the following anonymized GitHub repository: https://anonymous.4open.science/r/tas-revisions-7650.
@@
Figure S2: Trajectory reconstruction and prediction by best-case-per-$\lambda$ PINN and MAGI for the SEIR model in the fully observed case; for this SEIR example, the practitioner-style implementation (van Herten et al. 2022) outperformed the DeepXDE implementation across all $\lambda$ values. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Each solid blue curve is the estimate from one dataset. Top row: MAGI estimates. Lower five rows: best-case-per-$\lambda$ PINN estimates over all five tested $\lambda$ hyperparameter values.
Figure S3: PINN results from the DeepXDE implementation (Lu et al. 2021), which performs worse than the practitioner-style implementation (van Herten et al. 2022) used in the main text for this SEIR example. Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets in the fully observed case. Lower values indicate better performance. The y-axis is displayed on a logarithmic scale for improved visualization. Top row: in-sample trajectory reconstruction (Equation (9)); bottom row: future forecasting (Equation (10)). The three columns correspond to the three system components $E$, II, and $R$. In each panel, the leftmost boxplot corresponds to MAGI, and the remaining boxplots correspond to best-case-per-$\lambda$ DeepXDE PINN under different hyperparameter values $\lambda$. The dashed vertical line separates MAGI and PINN results.
Figure S4: PINN results from the DeepXDE implementation (Lu et al. 2021), which yields worse parameter estimation performance than the practitioner-style implementation (van Herten et al. 2022) used in the main text for this SEIR example. Boxplots showing absolute parameter estimation errors across 100 datasets in the fully observed case. Lower values indicate better performance. The y-axis is displayed on a logarithmic scale for improved visualization. Top row: errors for the original SEIR parameters $\beta$, $\gamma$, and $\sigma$. Bottom row: errors for $R_{0}$, peak timing, and peak intensity—our three quantities of interest. In each panel, the leftmost boxplot corresponds to MAGI, and the remaining boxplots correspond to best-case-per-$\lambda$ DeepXDE PINN under different hyperparameter values $\lambda$. The dashed vertical line separates MAGI and PINN results.
@@
Figure S5: Trajectory reconstruction and prediction by best-case-per-$\lambda$ PINN and MAGI for the SEIR model in the missing component case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Each solid blue curve is the estimate from one dataset. Top row: MAGI estimates. Lower five rows: best-case-per-$\lambda$ PINN estimates over all five tested $\lambda$ hyperparameter values.
Figure S6: PINN results from the DeepXDE implementation (Lu et al. 2021) for the missing $E$ component case; as in the fully observed case, the DeepXDE-based PINN performs worse than the practitioner-style implementation used in the main text. Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets. The legend and layout are identical to Figure 5 in the main text; see the caption there for details. In each panel, the leftmost boxplot corresponds to MAGI, and the remaining boxplots correspond to best-case-per-$\lambda$ DeepXDE PINN under different hyperparameter values $\lambda$. The dashed vertical line separates MAGI and PINN results.
Figure S7: PINN results from the DeepXDE implementation (Lu et al. 2021) for parameter estimation in the missing $E$ component case; consistent with the trajectory results, the DeepXDE-based PINN yields worse performance than the practitioner-style implementation used in the main text. Boxplots showing absolute parameter estimation errors across 100 datasets. The legend and layout are identical to Figure 6 in the main text; see the caption there for details. In each panel, the leftmost boxplot corresponds to MAGI, and the remaining boxplots correspond to best-case-per-$\lambda$ DeepXDE PINN under different hyperparameter values $\lambda$. The dashed vertical line separates MAGI and PINN results.
S5 Additional Lorenz chaotic (butterfly) results